Consider the compound interest equation B(t)=100(1.1664)t

Assume that n=2, and rewrite B(t) in the form B(t)=P(1+r/n)nt


What is the interest rate, r, written as a percentage?

Enter your answer as a whole number, like this: 42


1
Expert's answer
2021-05-28T10:03:07-0400

Solution:

"B(t)=P(1+\frac rn)^{nt}"

"B(t)=100(1.1664)^t=100(1+0.1664)^t"

Put "n=2"

"B(t)=100(1+\frac{0.1664}{2})^{2t}"

Here, interest rate, "r=0.1664=16.64\%\approx17\%"

Answer is 17.